Given that 192 = 2^x multiplied by 3^y find: A = x + y
Answer by jsmallt9(3758)

The problem tells us that 192 is some power of 2 times some power of 3. TO find out what these powers are (and what they add up to), let's factor 192 into 2's and 3's:
192 = 2*96 = 2*2*48 = 2*2*2*24 = 2*2*2*2*12 = 2*2*2*2*2*6 = 2*2*2*2*2*2*3 = 2^6 * 3^1.
So x = 6, y = 1 and x+y = 7